OpenStudy (akgorgeous):

REVIEWING FOR MY FINAL!! HELP NEEDED- Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 8x and y = 2^x + 2 intersect are the solutions of the equation 8x = 2^x + 2. Part B: Make tables to find the solution to 8x = 2^x + 2. Take the integer values of x between −3 and 3. Part C: How can you solve the equation 8x = 2^x + 2 graphically?

OpenStudy (akgorgeous):

@wanell

OpenStudy (wanell):

if they y values are the same then you can equate these 2 equations y = 8x and y = 2x + 2 8x=2x+2

OpenStudy (wanell):

Because they're both " y = something", you can substitute them into each other. y=8x and y=2x+2 8x=2x+2

OpenStudy (wanell):

for part A: y=8x y=2x+2 for part B 8x=2x+2 for part C: 8x=2x+2

OpenStudy (wanell):

As for part B, you can make a table, first make a function. Like last time, remember that y = f(x) and f(x) = y. That being said, y = 8x ==> f(x) = 8x, now make a table with x values of -3 to 3. when x = 2, f(2) = 8(2) = 16 therefore f(2) = 16 for the first function. The point in this is to see when they both have the same y value, when they do, they're both equal to each other. That's when they intersect.

OpenStudy (wanell):